Assume that X has a uniform distribution over [0; 1] and that Y has the uniform distribution over


[2; 3]: Which of the following statements are true and which are false? Justify your answers!


(a) P (X < Y) = 1.


(b) Since X is smaller than Y , P (X < 1) > P (Y < 1).


(c) There are some values a for which P (X < a) = P (Y < a).

1
Expert's answer
2022-05-11T11:02:38-0400

(a) The first statement is correct. Suppose that "X,Y:\Omega\rightarrow\mathbb{R}". Denote by "\Omega_1\sub\Omega" a subset of "\Omega" such that "X(w)\notin[0,1]" for all "w\in\Omega_1". Denote by "\Omega_2\sub\Omega" a subset of "\Omega" such that "Y(w)\notin[2,3]", for all "w\in\Omega_2". Denote by "\Omega_3" a subset of "\Omega" satisfying "X(w)<Y(w)" for all "w\in\Omega_3". It is clear that "\Omega_3\sub\Omega_1\cup\Omega_2". But "P(\Omega_1\cup\Omega_2)\leq P(\Omega_1)+P(\Omega_2)=0". We receive that "P(X<Y)=P(\Omega_3)=0"

(b) "P(Y<1)=0", since "Y" has the uniform distribution over "[2,3]". "P(X<1)=1". We receive: "P(X<1)>P(Y<1)."

(c) We can take "a=0". We get "0=P(X<0)=P(Y<0)". due to the definition of the uniform distribution.